# CERN-TH-2024-154

### v2

@@ I Introduction
In the present context, a more economical model is to instead use the dark Higgs since it is already present to generate the $A^{\prime}$ mass. Ref. Dror:2018pdh showed that if $h$ is displaced from the minimum of its potential at the end of inflation, then $A^{\prime}$ can obtain the desired relic density by parametric resonant production. That study assumed the initial value of $h$ could be arbitrary. Here we note that stochastic growth of fluctuations during inflation gives rise to a calculable estimate for this starting value, in terms of the inflationary energy scale Chatrchyan:2022dpy . Moreover, simple perturbative decays $h\to A^{\prime}A^{\prime}$ are sufficient to give the desired relic density. Overproduction of $A^{\prime}$ leads to new constraints on $\epsilon$; this is discussed in Section III.
@@
Figure 1: Cyan: excluded region from stellar cooling plus perturbative unitarity of dark Higgs self-coupling, derived in Section II. Magenta: region excluded by overproduction of $A^{\prime}$ from the decays of inflationary misaligned dark Higgs, derived in Section III. Red: Excluded region due to overproduction of $A^{\prime}$ via freeze-in, derived in Section III. Grey: Combination of astrophysical, cosmological and laboratory constraints on dark photons, adapted from Ref. Caputo:2021eaa . Green: Excluded region by cosmic string formation, from Long:2019lwl .
@@ II Higgsed Dark photon constraints
If a dark Higgs $h$ gives rise to the dark photon mass by getting a VEV $v$, it can participate in processes that would otherwise only involve $A^{\prime}$ Ahlers:2008qc . The effect is important when $m_{A^{\prime}}$ is negligible compared to the energy scales $T$ of interest. Then one can work in the basis where $h$ acquires a millicharge and can thereby be emitted from stars, leading to anomalous cooling by electromagnetic processes. To avoid ensuing bounds Davidson:2000hf , one must assume that $m_{h}>T$ so that the emission is kinematically blocked.
However, there is a theoretical upper limit on $m_{h}^{2}=2\lambda v^{2}$, when written in terms of the VEV and self-coupling $\lambda$ in the potential $V=\lambda(|H|^{2}-v^{2}/2)^{2}$, where $H$ represents the complex field. Perturbative unitarity of $\lambda$ implies that $\lambda<8\pi/3$, while $v$ is related to the dark photon mass by $m_{A^{\prime}}=gv$. These can be tied together by making a generic assumption about the size of the kinetic mixing, in the absence of any special cancellations: the typical contribution from a loop of a heavy particle carrying both U(1) charges gives
@@
Then we can eliminate $v,\,g,\,m_{h}$ in favor of $m_{A^{\prime}},\,\epsilon,\,T$,
@@
in order to avoid $h$ emission by stars of temperature $T\sim 1$ keV. This bound only applies in the region of parameter space $\epsilon\gtrsim 10^{-14}$ where it can soften the limit that was obtained for massless $h$, resulting in the cyan curve shown in Fig. 1.
@@ III Inflationary generation of $A^{\prime}$ via dark Higgs
Ref. Dror:2018pdh showed how parametric resonance of an oscillating dark Higgs $h$ could efficiently produce $A^{\prime}$ dark matter. Their analysis relied upon arbitrary starting values of $h$ away from its minimum. We note that a definite initial value can be predicted from the stochastic misalignment of $h$ that gets generated by its fluctuations during inflation. Furthermore we employ the simpler mechanism of perturbative decays $h\to A^{\prime}A^{\prime}$ to produce the dark matter. We find that this works in the regime where kinetic mixing is sufficiently small so that $A^{\prime}$ is long-lived compared to the age of the Universe. On the other hand, $h\to A^{\prime}A^{\prime}$ decays are unsuppressed for small $\epsilon$ or $g$; in the limit $g\to 0$, the width goes as $\Gamma\simeq g^{2}m_{h}^{3}/(8\pi m_{A^{\prime}}^{2})=\lambda m_{h}/(4\pi)$ due to the enhanced coupling to longitudinal $A^{\prime}$s Araki:2020wkq . For the parameters of interest, this leads to decays well before big bang nucleosynthesis, except in a small corner (upper left in Fig. 1) that is already excluded by other constraints.
@@
This leads to the constraint $\epsilon\gtrsim 1.4\times 10^{-12}$, which is shown in Fig. 2 (magenta region).
@@ IV Freeze-out by $s$-channel $A^{\prime}$ exchange
by combining Eqs. (11-12) with the working assumption $m_{A^{\prime}}>m_{\chi}$. Since thermal freezeout requires $m_{\chi}\gtrsim 1\,$MeV, it only applies for heavier $A^{\prime}$ than depicted in Fig. 1. In fact, we will show that $m_{A^{\prime}}$ should exceed $\sim 50$ MeV. The excluded region is shown in Fig. 2.
@@
which is much larger than the typical annihilation cross section. Nevertheless, since one generically expects tat $\epsilon\ll g_{\chi}$, it is interesting to determine the minimum ratio $\epsilon/g_{\chi}$ compatible with the combined relic abundance and self-interaction constraints. Combining Eqs.  (11-15), we find that $\epsilon/g_{\chi}$ for thermally produced dark matter is bounded from below by
@@
[…]
@@ V Conclusions
We further derived new constraints on dark photon-mediated thermal dark matter, in the regime where its mass is less than $m_{A^{\prime}}$, using the one-loop estimate for $\epsilon$ and constraints on on self-interacting dark matter from the Bullet cluster. These considerations rule out $m_{A^{\prime}}<46\,$MeV, and place strong (model-independent) upper and (model-dependent) lower bounds on the $\epsilon$ that can diminish the allowed parameter space significantly.

### v3

@@ I Introduction
Figure 1: Cyan: excluded region from stellar cooling plus perturbative unitarity of dark Higgs self-coupling, derived in Section II. black: region excluded by overproduction of $A^{\prime}$ from the decays of inflationary misaligned dark Higgs, derived in Section III. Red: Excluded region due to overproduction of $A^{\prime}$ via freeze-in, derived in Section III. Orange: Region of parameter space where the decays $h\rightarrow A^{\prime}A^{\prime}$ yield relativistic dark photons today. Grey: Combination of astrophysical, cosmological and laboratory constraints on dark photons, adapted from Ref. Caputo:2021eaa . Green: Region excluded by cosmic string formation, from Ref. Long:2019lwl .
In the present context, a more economical model is to instead use the dark Higgs since it is already present to generate the $A^{\prime}$ mass. After shifting the dark Higgs $h$ by its VEV, $h\to h+v$, and using $v=m_{A^{\prime}}/g$ to eliminate $v$, the Lagrangian is
$\mathcal{L}=$ $-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{4}F^{\prime}_{\mu\nu}{F^{\prime}}^{\mu\nu}-\frac{\epsilon}{2}F_{\mu\nu}{F^{\prime}}^{\mu\nu}$
$+\frac{m_{A^{\prime}}^{2}}{2}A^{\prime}_{\mu}{A^{\prime}}^{\mu}+{A^{\prime}}^{\mu}g_{\chi}\bar{\chi}\gamma_{\mu}\chi$
$+gm_{A^{\prime}}hA^{\prime}_{\mu}{A^{\prime}}^{\mu}+\frac{1}{2}g^{2}h^{2}A^{\prime}_{\mu}{A^{\prime}}^{\mu}.$ (1)
where ${F^{\prime}}^{\mu\nu}=\partial^{\mu}{A^{\prime}}^{\nu}-\partial^{\nu}{A^{\prime}}^{\mu}$, and $\chi$ is a dark fermion. The term involving $\chi$ will only become relevant in Section IV, where the dark photon will be considered as a mediator. Ref. Dror:2018pdh showed that if $h$ is displaced from the minimum of its potential at the end of inflation, then $A^{\prime}$ can obtain the desired relic density by parametric resonant production. That study assumed the initial value of $h$ could be arbitrary. Here we note that stochastic growth of fluctuations during inflation gives rise to a calculable estimate for this starting value, in terms of the inflationary energy scale Chatrchyan:2022dpy . Moreover, simple perturbative decays $h\to A^{\prime}A^{\prime}$ are sufficient to give the desired relic density. Overproduction of $A^{\prime}$ leads to new constraints on $\epsilon$; this is discussed in Section III.
@@ II Higgsed Dark photon constraints
If a dark Higgs $h$ gives rise to the dark photon mass by getting a VEV $v$, it can participate in processes that would otherwise only involve $A^{\prime}$ Ahlers:2008qc . The effect is important when $m_{A^{\prime}}$ is negligible compared to the energy scales $T$ of interest. Then one can work in the basis where $h$ acquires a millicharge and can thereby be emitted from stars, leading to anomalous cooling by electromagnetic processes, mainly driven by the plasmon decay channel $\gamma^{\star}\rightarrow A^{\prime}+h$. To avoid ensuing bounds Davidson:2000hf , one must assume that $m_{h}>T$ so that the emission is kinematically blocked.
However, there is a theoretical upper limit on $m_{h}^{2}=2\lambda v^{2}$, when written in terms of the VEV and self-coupling $\lambda$ in the potential $V=\lambda(|H|^{2}-v^{2}/2)^{2}$, where $H$ represents the complex field. Perturbative unitarity of $\lambda$ implies that $\lambda<8\pi/3$, while $v$ is related to the dark photon mass by $m_{A^{\prime}}=gv$. These can be tied together by making a generic assumption about the size of the kinetic mixing, in the absence of any special cancellations: the typical contribution from a loop of a heavy particle $\phi$ carrying both U(1) charges gives Pich:1998xt
$\epsilon=\frac{ge}{16\pi^{2}}\mathrm{log}\Big{(}\frac{m_{\phi}^{2}}{\mu_{0}^{2}}\Big{)}$ (2)
where $\mu_{0}$ is the renormalization scale.
The logarithm could accidentally be small, or there could be a cancellation mechanism, such as two nearly equal-mass particles with opposite charges contributing in the loop. In such cases we do not gain any new constraints from the consideration of kinetic mixing. In the present work we are interested in the generic situation in which the log is of order unity, and we henceforth set it to 1.
Then we can relate the kinetic mixing $\epsilon$ with the gauge coupling $g$
@@
Further eliminating $v,\,g,\,m_{h}$ in favor of $m_{A^{\prime}},\,\epsilon,\,T$, we obtain
@@
in order to avoid $h$ emission by stars of temperature $T\sim 1$ keV. This bound applies in the region of parameter space where it can soften a previously derived limit for sufficiently light $h$ An:2013yua . That reference obtained constraints of order $\epsilon\lesssim 10^{-13}$ from the anomalous cooling of horizontal branch stars, but this relied on the ad hoc assumption that $g=0.1$. Refined stellar cooling constraints from the tip of the red-giant branch lie on the same order of magnitude Fung:2023euv . When relating $g$ and $\epsilon$ via Eq. 3, we find instead the limit
$\epsilon\lesssim\sqrt{\frac{8\times 10^{-15}e}{16\pi^{2}}}\lesssim 3.9\times 10^{-9}.$ (5)
The scaling of the bound in $\epsilon$ with $g$ arises from the plasmon decay rate governing the energy loss in stars, which scales as $\epsilon^{2}g^{2}$ Raffelt:1996wa . Our combined stellar cooling constraints are shown in Fig. 1 as the cyan region.
@@ III Inflationary generation of $A^{\prime}$ via dark Higgs
Ref. Dror:2018pdh showed how parametric resonance of an oscillating dark Higgs $h$ could efficiently produce $A^{\prime}$ dark matter. Their analysis relied upon arbitrary starting values of $h$ away from its minimum. We note that a definite initial value can be predicted from the stochastic misalignment of $h$ that gets generated by its fluctuations during inflation. Furthermore we employ the simpler mechanism of perturbative decays $h\to A^{\prime}A^{\prime}$ to produce the dark matter. We find that this works in the regime where kinetic mixing is sufficiently small so that $A^{\prime}$ is long-lived compared to the age of the Universe. On the other hand, $h\to A^{\prime}A^{\prime}$ decays are unsuppressed for small $\epsilon$ or $g$; in the limit $g\to 0$, the width goes as $\Gamma\simeq g^{2}m_{h}^{3}/(8\pi m_{A^{\prime}}^{2})=\lambda m_{h}/(4\pi)$ due to the enhanced coupling to longitudinal $A^{\prime}$s Araki:2020wkq . For the parameters of interest, this leads to decays well before big bang nucleosynthesis, except in the orange shaded region in Fig. 1, which is mostly excluded by other constraints. We note that a small triangular region remains viable, around $m_{A^{\prime}}\sim 10^{-17}-10^{-14}$ eV and $\epsilon\sim 10^{-12}-10^{-8}$.
@@
This leads to the constraint $\epsilon\gtrsim 1.4\times 10^{-12}$, which is shown in Fig. 2 (black region).
@@ IV Freeze-out by $s$-channel $A^{\prime}$ exchange
by combining Eqs. (14-15) with the working assumption $m_{A^{\prime}}>m_{\chi}$. Since thermal freezeout requires $m_{\chi}\gtrsim 1\,$MeV, it only applies for heavier $A^{\prime}$ than depicted in Fig. 1. In fact, we will show that $m_{A^{\prime}}$ should exceed $\sim 50$ MeV. The excluded region is shown in Fig. 2.
@@
which is much larger than the typical annihilation cross section. Nevertheless, since one generically expects tat $\epsilon\ll g_{\chi}$, it is interesting to determine the minimum ratio $\epsilon/g_{\chi}$ compatible with the combined relic abundance and self-interaction constraints. Combining Eqs.  (14-18), we find that $\epsilon/g_{\chi}$ for thermally produced dark matter is bounded from below by
@@
In the preceding, we assumed that $g_{\chi}=g=16\pi^{2}\epsilon/e$. Conceivably $g_{\chi}$ could be much larger than $g$. Considering representative values $g_{\chi}=0.1\,e,\,e,\,10\,e$, and mass ratio $m_{\chi}/m_{A^{\prime}}=1/3$, we find interesting lower bounds on $\epsilon$ versus $m_{A^{\prime}}$ illustrated in Fig. 2 (right). This probes a region of parameter space which has not yet been excluded by collider or beam-dump experiments.
@@ V Conclusions
We further derived new constraints on dark photon-mediated thermal dark matter, in the regime where its mass is less than $m_{A^{\prime}}$, using the one-loop estimate for $\epsilon$ and constraints on self-interacting dark matter from the Bullet cluster. These considerations rule out $m_{A^{\prime}}<46\,$MeV, and place strong (model-independent) upper and (model-dependent) lower bounds on the $\epsilon$ that can diminish the allowed parameter space significantly.